Jon recently drove to visit his parents who live 225 miles away. On his way there his average speed was 24 miles per hour faster than on his way home. If Jon spent a total of 10 hours driving, find the two rates.
Answer by mananth(16949)

LET HIS SPEED ON HIS WAY HOME BE X MPH
Time taken in the return journey = 225/x
His speed while going was x+24
time taken in forward journey 225/ x+24
225/x + 225 / x+24 = 10 hours
225(x+24)+225x= 10*x(x+24)
225x+5400 +225x= 10x^2 +240x
10x^2-210x-5400=0
x^2-21x-540=0
x^2-36x+15x-540=0
x(x-36)+15(x-36)=0
(x-36)(x+15)=0
x= 36 mph on way home
while going it was 36+24 = 60 mph
Ananth